This is just a brief update on misc. activities over the last 2 month in preparations for the upcoming RHIC run. I hope this is reasonable complete, if something has been forgotten excuse me. - DC modules t3 was repaired. (pawel,radek,bob/john,ed) - Si has been reconfigured for enhanced flow-measurements *(Steve,Bjoern,Hiro) {I hope we can have a brief description , as a note what the present geometry is) - TPC's has been checked out (readout, gasmixture adjusted for proper Vdrift using monitors) (Ramiro) - T1 and T2 are moved closer to shields, to have active volume slight farther way from edges. (Ramiro,Ed) - T1,T2 and T3 surveyed (CA-D; waiting for results) - BB system chekced out (Burt) - DAQ and FEH computersystem reconfigured with new machines, different network-config -the RUNDB, and web server is now on brahms-rundb0. and brahms-web respectively. The db server is functional, online scripts changed to used these, the web server will be completely functional very soon. (Hiro, Bob) HV systems have been checked out, (some MF failures on old systems). Filters, fans, NIM, camac checked out (some still have problems) (Bob, John, Hiro) The RICH will be refilled just before the run (even though it has held pressure quite well).(Ramiro) regards Flemming ---------------------------------------------------------------- Flemming Videbaek Physics Department Brookhaven National Laboratory e-mail: videbaek@bnl.gov phone: 631-344-4106 _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Fri Nov 26 11:54:27 2004
This archive was generated by hypermail 2.1.8 : Fri Nov 26 2004 - 11:54:52 EST